OpenStudy (xapproachesinfinity):

how do i factor x^2-x-6

OpenStudy (jojo4eva):

(x-3)(x+2)

OpenStudy (jojo4eva):

you find factors of -6 that would equal -x which would be -3 and 2

OpenStudy (xapproachesinfinity):

@ganeshie8

OpenStudy (anonymous):

You can use this calculator -> http://www.algebrahelp.com/calculators/expression/factoring/calc.do?expression=-12x+-+24

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{x^2-x-6\\ =x^2-2\times \dfrac{x}{2}+\dfrac{1}{4}-\dfrac{1}{4}-6\\ =(x-\dfrac{1}{2})^2-\dfrac{25}{4}\\ =(x-\dfrac{1}{2})^2-(\dfrac{5}{2})^2\\ =(x-\dfrac{1}{2}-\dfrac{5}{4})(x-\dfrac{1}{2}+\dfrac{5}{4})}\end{align}\)

OpenStudy (jojo4eva):

@helpme25

OpenStudy (xapproachesinfinity):

I posted this for my class speech about Open Study lol thanks

OpenStudy (mathmath333):

oh which grade u study

OpenStudy (xapproachesinfinity):

I'm college math undergraduate